# references.bib Weak Ties Explain Open Source Innovation

### v1

@@ V Discussion
Next we discuss our results in the broader context of the literature and highlight budding directions for future research.
V-A Summary of Main Results
@@
Established social science theory describes the mechanisms through which weak ties in a social network can be particularly “strong”—they often help bridge disjoint parts of the network, thus providing access to diverse and novel information, which in turn can be used towards better outcomes. Our study provides the first empirical link (that we know of) between this theoretical mechanism and a practical and important software engineering outcome—the emergence of innovation in a software project. Our hypotheses assert that the many user-to-user and user-to-artifact interactions possible on the GitHub platform, ranging from high-effort ones like making changes to a codebase to relatively trivial ones like starring a repository, result in network ties of varying strength, along which knowledge can flow. And, in particular, that weak ties are instrumental for this knowledge flow, as the GeoNotebook anecdote in the Introduction (Figure 1) would suggest.
@@
Using a sophisticated methodology to reconstruct interaction networks, estimate tie strength, compute network informational diversity, and estimate project innovativeness, coupled with a robust statistical analysis, we found clear evidence supporting two of our three hypotheses: it’s not the amount (H), but rather both the diversity of information available to a focal project’s core developers (H), as well as the extent to which this information diversity comes from weak ties (H), that explain some of the variance in how novel (atypical) the combination of software packages eventually imported in the focal project are.
V-B Implications
@@
One way to interpret our results is that lurking on the GitHub platform (where we see star gazing as lurking rather than actively contributing) has quantifiable benefits. If indeed people draw inspiration from things they’ve seen and apply those ideas to something they will build next,555Note, our study does not have sufficient power to make causal claims. platform designers should consider facilitating this process more. For example, GitHub has a way of highlighting “what the community is most excited about today” on its Trending page. It could be beneficial if such algorithms that determine what’s trending (allegedly666https://medium.com/@DiggerHQ/how-we-got-into-github-trending-c281f3b06df9 currently based on star count growth rate) also take into account the overall informational diversity of the highlighted repositories, or even better, the contribution to enhancing the informational diversity for a given user, i.e., if the recommendations were personalized.
In addition, research has shown that developers tend to join, or pay attention to, projects that are technically familiar to them [dey2021representation, fang2023matching], and automated project recommendation tools have been proposed, that recommend projects to developers with related technical background [zhang2017devrec]. Our results suggest that such approaches maximizing technical similarity might be counterproductive, at least in terms of enabling innovation.
@@
Our results also raise questions about the GitHub culture of (implicitly) rewarding code contributions the most, since one’s commit history tends to be most highly visible in the platform. Much has been written about how other developers [marlow2013impression] and even recruiters [capiluppi2012assessing] use such signals. In contrast, we are finding evidence that well-informed but not necessarily highly active developers may also be experts at their craft, at least insofar as the novelty of their output.
@@
It’s also worth thinking about what tracking and giving credit to ideas, not just code artifacts, might look like in this domain. In scientific research, authors are expected to acknowledge (and cite) prior work that influenced their thinking. In turn, these citation networks encode invaluable information about scientific progress and scientific collaboration [fortunato2018science]. What might such a system look like in software development? Anecdotally, it seems like developers already use many ad-hoc ways to credit ideas, e.g., including links in source code comments [hata20199] and even referring to research papers in repository README files [wattanakriengkrai2022github] or code [inokuchi2019academia]. But we still don’t understand how common the practice is, how complete the mentions are, and what the supply chains of ideas look like.
@@
Questions around the value of having access to diverse information have also come up in research focused on demographic team composition, e.g., around gender [vasilescu2015gender] and race/ethnicity [shameer2023relationship]. This work posits similar underlying mechanisms (surface-level demographic variables act as a proxy for deeper-level, less observable differences in backgrounds, skills, approaches to solve problems, etc) to the one we measure here more directly. Now that such measurements are becoming possible, it would be interesting to further test the relationships between team-level surface attributes, informational or network diversity, and outcomes.
@@
Finally, we can’t help but wonder how the proliferation of AI-based assistants will impact the innovativeness of the software being created. A large language model has certainly seen more diverse information in its training than any individual developer. But is it trained to expose that diversity, e.g., when generating code snippets automatically, or might AI-generated code end up looking more like a regression to the mean?
VI Conclusions
@@
Much like in Granovetter’s jobs search study we started from in the Introduction, wherein most often people found their jobs through weak-tie acquaintances rather than strong-tie friends or family, we conclude that open-source developers may have found useful ideas to help them create innovative projects through passive GitHub interactions, like star gazing, more often than through active ones like committing to a repository.
VII Data Availability
@@
Our replication package is available online777https://github.com/icsesubmission/replication_package_icse25_novelty and it includes aggregated data in CSV format and R scripts that reproduce the tables and figures in the paper.

### v2

@@ 5. Discussion
Next we discuss our results in the broader context of the literature and highlight budding directions for future research and practice.
@@
Summary of Main Results. Established social science theory describes the mechanisms through which social network weak ties can be particularly “strong” – they can bridge disjoint parts of the network, providing access to diverse information, which in turn can be used towards better outcomes. Our study provides the first empirical link (that we know of) between this theoretical mechanism and an important software engineering outcome – the emergence of innovation in a software project. We hypothesized that the many user-to-user and user-to-artifact interactions possible on the GitHub platform, ranging from high-effort ones like making changes to a codebase to relatively trivial ones like starring a repository, result in network ties of varying strength, along which knowledge can flow. And, in particular, that weak ties are instrumental for this knowledge flow, as the GeoNotebook anecdote in the Introduction (Figure 1) would suggest.
@@
Using a sophisticated methodology to reconstruct interaction networks, estimate tie strength, compute network informational diversity, and estimate project innovativeness, coupled with a robust statistical analysis, we found clear evidence (albeit only correlational) supporting two of our three hypotheses: on average, it’s not the amount (H), but rather both the diversity of information available to a focal project’s core developers (H), as well as the extent to which this information diversity comes from weak ties (H), that explain some of the variance in how novel (atypical) the combination of software packages imported in a focal project is.
Effect Sizes. Some readers may question whether our main variables explain enough variance, but small effect sizes are exactly what we should expect when studying innovation in complex socio-technical systems. Software development operates in an extraordinarily high-dimensional space where countless factors influence outcomes, making any single mechanism necessarily modest in isolation. The theoretical mechanism of weak ties operates through subtle informational advantages that accumulate over time rather than dramatic singular influences, and these effects must compete against an already information-rich environment of documentation, tutorials, and online resources. Moreover, as discussed in the beginning of our paper, the weak ties theory has decades of robust empirical support across diverse domains, consistently showing modest but meaningful effects that compound over time. We’re far from the first to discover this effect – at best, we’re the first to offer some supporting evidence from the open-source software development context. In addition, most software projects engage in predictable recombination of existing packages, with genuinely novel combinations occurring primarily in the tail of the distribution where even small probability shifts can have outsized impacts. The consistency of effects across multiple strata, coupled with statistical significance despite the noisy environment, actually strengthens confidence in the underlying mechanism – large effect sizes in network research are often suspicious because they suggest implausibly crude influences that overwhelm all other factors, which rarely occurs in real complex systems. Next, we discuss the implications of our results.
@@
Platform Design. One way to interpret the strength-of-weak-ties theory in our context is that lurking on the GitHub platform (i.e., we see starring repositories as lurking rather than actively contributing) has quantifiable benefits. If indeed people draw inspiration from things they’ve starred and apply those ideas to something they will build next, as the theory would suggest and our quantitative results support, platform designers should consider facilitating this process more. For example, GitHub has a way of highlighting “what the community is most excited about today” on its Trending page. It could be beneficial if this algorithm that determines what is trending (allegedly based on the star growth rate) also took into account and increased the overall informational diversity of the highlighted repositories (as computed, e.g., using an embedding-based approach like ours). Better still, the recommendations could be personalized, i.e., the algorithm could highlight repositories that enhance information diversity for a given user.
@@
Our results also raise questions about the GitHub culture of (implicitly) rewarding code contributions the most, since one’s commit history tends to be highly visible in the platform. Much has been written about how other developers (Marlow et al., 2013) and even recruiters (Capiluppi et al., 2012) use such signals. In contrast, we are finding evidence that well-informed but not necessarily highly active developers may also be experts at their craft, at least insofar as the novelty of their output. It’s worth thinking about how to feature (and reward) such lurking more prominently in the platform design.
@@
It’s also worth thinking about what tracking and giving credit to ideas, not just code artifacts, might look like in this domain. In scientific research, authors are expected to cite prior work that influenced their thinking. In turn, these citations form networks encoding invaluable information about scientific progress and scientific collaboration (Fortunato et al., 2018). What might such a system look like in software development? Anecdotally, it seems like developers already use many ad-hoc ways to credit ideas, e.g., including links in source code comments (Hata et al., 2019) and even referring to research papers in repository README files (Wattanakriengkrai et al., 2022) or source code (Inokuchi et al., 2019). But we still don’t understand how common the practice is, how complete the mentions are, and what the supply chains of ideas look like. Nor do we have ways of systematically tracking ideas as part of the platform design, which is worth exploring.
Theory. More broadly, there are many opportunities for developers for relatively low-effort interaction with each other and with new technology besides the one we measured here (starring). These include using social media to stay current (Singer et al., 2014; Fang et al., 2022; Wyrich and Bogner, 2024), listening to podcasts (Engzell and Norrman, 2023), creating or consuming video content online (Chattopadhyay et al., 2021), and participating in developer conferences (Truong et al., 2022) or trainings. Since our study is not exploratory (bottom-up), but rather tests hypotheses drawn from a much more general, widely-validated theory (top-down), it’s reasonable to expect, given the theory, that the same “strong weak ties” mechanism could also apply in these scenarios. The famous “water-cooler effect,” that advocates of in-person work environments often quote, is another example of the same mechanism – casual encounters, typically with weak ties (work acquaintances rather than close colleagues), can help exchange new ideas that spark creativity (Brucks and Levav, 2022). We need more research to empirically investigate these and other types of weak ties and their innovation-enabling potential, as it would be important not only to collect more evidence in support of the mechanism, but also to rank the different interactions in terms of their effectiveness. For example, many technology companies support their engineers regularly attending conferences and these experiences are likely beneficial. Could we quantify these benefits both in general, and relative to, say, consuming technical social media content?
Separately, our work aligns well with Baltes and Diehl’s (Baltes and Diehl, 2018) software development expertise theory. For example, our lurking can be considered as an example of behavior that contributes to a developer’s knowledge base, or as an example of continuous learning; similarly, the GitHub platform is the work context that influences expertise development, by facilitating exposure to diverse projects. But perhaps more interestingly, while the theory emphasizes deliberate practice for expertise development, our work offers a complementary perspective and an opportunity for theory refinement: casual observation of diverse projects may contribute to innovative thinking in ways that focused practice alone might not.
Exploration vs Exploitation. Research has shown that developers are more likely to join projects that are technically more familiar to them (Dey et al., 2021; Fang et al., 2023). Various automated project recommendation tools have also been proposed, that match projects to developers with the closest technical background (Zhang et al., 2017). Such matching is likely useful, as open-source projects depend heavily on contributions from volunteers, and contributors can be in short supply (Coelho and Valente, 2017). At the same time, our results suggest that maximizing technical similarity between a project and the backgrounds of its contributors might be counterproductive in terms of enabling innovation. More research is needed to explore this potential trade-off between exploration (when divergent thinking is likely beneficial) and exploitation (when narrow focus is required) (Liu et al., 2021) in a software development context, particularly open source. For example, while tracking novelty computationally is becoming increasingly feasible (our work demonstrates this), we are still lacking in our understanding of when in the lifecycle of an open-source project innovations occur and who is responsible for them. Future work could start exploring how novel design emerges at the onset of a project and how that compares to, e.g., feature request discussions (by definition opportunities for innovation) occurring throughout a project’s lifetime.
@@
Diversity. Questions around the value of having access to diverse information have also come up in research focused on demographic team composition, e.g., around gender (Vasilescu et al., 2015) and race / ethnicity (Shameer et al., 2023). This past work posits a similar underlying mechanism (surface-level demographic variables act as a proxy for deeper-level, less observable differences in backgrounds, skills, approaches to solve problems, etc) to the one we measure here more directly. Now that such measurements are becoming possible, it would be interesting to further test the relationships between team surface-level attributes, informational or network diversity, and outcomes.
@@
AI and Creativity. Finally, we can’t help but join colleagues (Jackson et al., 2024) in wondering how AI-based assistants will impact the innovativeness of the software being created. A large language model has certainly seen more diverse information in its training than any individual developer. But is the user experience designed to expose that diversity, e.g., when generating code snippets automatically, or might AI-generated code end up looking more like a regression to the mean with current iteraction modalities?
6. Conclusion
@@
Much like in Granovetter’s jobs search study exposing the strength of weak ties, we conclude that open-source developers may also find useful ideas to help them create innovative projects through passive GitHub interactions, like starring repositories, more often than through active engagement like committing to a repository.
@@
Data Availability. Our replication package666https://github.com/icsesubmission/replication_package_icse26_novelty (Zenodo post acceptance) includes data and scripts to reproduce our tables and figures.
Appendix A Using transitivity to validate the ‘strength’ of different networks
In Section 3.3, we reason that interactions via commits, issues, and stars reflect ties of decreasing strength, and here we provide an empirical validation of this assumption through the comparison of transitivity among the three networks.
Transitivity measures the density of local connections in the local network (Newman, 2003). Consider the illustration in Figure 6: if node is connected to and , a triad is formed. In social network theory, if ’s connections to and are both strong ties, the triad is unstable and will likely evolve into a triangle where , , and are all connected. This evolution occurs because and have more opportunities to interact and increase their mutual familiarity due to their strong ties with . Thus, networks characterized by strong ties typically exhibit a greater number of triangles, as opposed to triads, compared to those with weak ties (Granovetter, 1973). Formally, this can be computed using the measurement of transitivity, defined as . A higher transitivity value (number of triangles relative to triads) indicates a network of stronger ties.
Similarly, we compare the transitivity values of our three networks, summarized in Table 4.777Since network transitivity is usually defined for undirected networks, we first convert our three networks (which are all directed) to their undirected form. That is, there is an undirected edge between nodes and if there is either an edge from to or from to . In addition, the reported number of nodes excludes isolates (i.e., nodes without edges) from each network, and the number of edges is calculated in the undirected version of each network. Overall, we observe approximately an order of magnitude () difference in transitivity values between each pair of networks. Specifically, the commit network displays the highest levels of transitivity, followed by the issue network, while the star network exhibits the lowest level. Thus, these findings are consistent with our theoretical understanding of tie strength.
| Interaction | #Nodes | #Edges | Transitivity () |
| Commits | |||
| Issues | |||
| Stars |
Appendix B Replication with the atypicality measure by Fang et al
We further test the robustness of our regression results by replicating the analysis presented in Table 3 using the original measurement of atypicality defined by Fang et al. (2024) as the outcome variable. The results of this robustness analysis are reported in Table 5. It is important to note that the range of the atypicality score in the main analysis differs from that in the robustness analysis. As a result, the magnitudes of the estimated coefficients may vary; however, the overall qualitative patterns remain consistent.
. Model I Model II Model III Model IV Variables of interest (H) 0.055∗∗∗ 0.009∗ 0.017∗∗∗ (0.0015) (0.0044) (0.0045) 0.009∗∗∗ 0.012∗ 0.020∗∗∗ (0.0022) (0.0054) (0.0059) (H) 0.042∗∗∗ 0.048∗∗∗ (0.0042) (0.0044) (H) 0.045∗∗∗ 0.053∗∗∗ (0.0056) (0.0060) Controls Org_owned 0.125∗∗∗ 0.071∗∗∗ 0.074∗∗∗ 0.076∗∗∗ (0.0058) (0.0118) (0.0116) (0.0118) (log) 0.016∗∗∗ 0.011∗∗∗ 0.010∗∗∗ 0.013∗∗∗ (0.0020) (0.0028) (0.0027) (0.0028) (log) 0.189∗∗∗ 0.093∗∗∗ 0.087∗∗∗ 0.094∗∗∗ (0.0053) (0.0114) (0.0112) (0.0114) (log) 0.236∗∗∗ 0.150∗∗∗ 0.153∗∗∗ 0.154∗∗∗ (0.0031) (0.0080) (0.0079) (0.0080) Fixed effect ✓ ✓ ✓ ✓ Observations 589,999 37,451 37,451 37,451 Adjusted R2 0.044 0.034 0.038 0.038 ∗p0.05; ∗∗p0.01; ∗∗∗p0.001
Appendix C Replication with Alternative Definitions of Core Developers and Interaction Periods
In Section 3.3, we constructed knowledge networks between projects based on interactions among their core developers within a 12-month period prior to their first contributions. To assess the robustness of our findings, we conducted replication analyses using alternative definitions of core developers and varying lengths of interaction periods, as reported in Table 3.
Following established practice (Mockus et al., 2002), we alternatively identify core developers as the top contributors to a project whose cumulative commits account for at least 80% of all project commits. The corresponding regression results using this alternative definition are presented in Table 6. Similarly, we extend the interaction window from 12 to 24 months prior to a core developer’s first contribution to the focal project when constructing the networks. The regression results based on this extended interaction period are presented in Table 7.
We also run the same model (as model IV in Table 3) with several other combinations of core developer identification approach and length of interaction periods, and the estimated coefficient of Variables of Interest are presented in Table 8
| Model I | Model II | Model III | Model IV | |
| Variables of interest | ||||
| (H) | 0.001∗∗∗ | 0.001 | 0.002∗∗∗ | |
| (0.0002) | (0.0006) | (0.0006) | ||
| 0.001∗∗ | 0.002∗ | 0.006∗∗∗ | ||
| (0.0002) | (0.0007) | (0.0008) | ||
| (H) | 0.007∗∗∗ | 0.008∗∗∗ | ||
| (0.0006) | (0.0006) | |||
| (H) | 0.004∗∗∗ | 0.006∗∗∗ | ||
| (0.0007) | (0.0008) | |||
| Controls | ||||
| Org_owned | 0.028∗∗∗ | 0.018∗∗∗ | 0.020∗∗∗ | 0.020∗∗∗ |
| (0.0006) | (0.0016) | (0.0015) | (0.0016) | |
| (log) | 0.002∗∗∗ | 0.001 | 0.0004 | 0.0007 |
| (0.0002) | (0.0004) | (0.0004) | (0.0004) | |
| (log) | 0.010∗∗∗ | 0.017∗∗∗ | 0.015∗∗∗ | 0.016∗∗∗ |
| (0.0005) | (0.0015) | (0.0015) | (0.0015) | |
| (log) | 0.043∗∗∗ | 0.025∗∗∗ | 0.025∗∗∗ | 0.026∗ |
| (0.0003) | (0.0010) | (0.0010) | (0.0010) | |
| Fixed effect | ||||
| Observations | 679,930 | 40,589 | 40,589 | 40,589 |
| Adjusted R2 | 0.053 | 0.064 | 0.068 | 0.070 |
| ∗p0.05; ∗∗p0.01; ∗∗∗p0.001 | ||||
| Model I | Model II | Model III | Model IV | |
| Variables of interest | ||||
| (H) | 0.002∗∗∗ | 0.0003 | 0.001∗∗ | |
| (0.0002) | (0.0005) | (0.0005) | ||
| 0.0002 | 0.002∗∗ | 0.007∗∗∗ | ||
| (0.0002) | (0.0006) | (0.0007) | ||
| (H) | 0.007∗∗∗ | 0.008∗∗∗ | ||
| (0.0004) | (0.0005) | |||
| (H) | 0.004∗∗∗ | 0.007∗∗∗ | ||
| (0.0007) | (0.0007) | |||
| Controls | ||||
| Org_owned | 0.027∗∗∗ | 0.020∗∗∗ | 0.022∗∗∗ | 0.021∗∗∗ |
| (0.0006) | (0.0014) | (0.0014) | (0.0014) | |
| (log) | 0.002∗∗∗ | 0.001∗ | 0.0010∗∗ | 0.0012∗∗∗ |
| (0.0002) | (0.0003) | (0.0003) | (0.0003) | |
| (log) | 0.013∗∗∗ | 0.015∗∗∗ | 0.015∗∗∗ | 0.015∗∗∗ |
| (0.0006) | (0.0014) | (0.0013) | (0.0013) | |
| (log) | 0.042∗∗∗ | 0.024∗∗∗ | 0.025∗∗∗ | 0.025∗∗∗ |
| (0.0003) | (0.0009) | (0.0009) | (0.0009) | |
| Fixed effect | ||||
| Observations | 587,710 | 51,571 | 51,571 | 51,571 |
| Adjusted R2 | 0.055 | 0.055 | 0.059 | 0.061 |
| ∗p0.05; ∗∗p0.01; ∗∗∗p0.001 | ||||
| Variables | |||||
Core
developer identifica -tion |
Interaction
period length (month) |
| Original | 6 | 0.001 | 0.005∗∗∗ | 0.006∗∗∗ | 0.006∗∗∗ |
| (0.0009) | (0.0011) | (0.0008) | (0.0008) | ||
| Alternative | 6 | 0.003∗∗ | 0.006∗∗∗ | 0.007∗∗∗ | 0.007∗∗∗ |
| (0.0008) | (0.0010) | (0.0008) | (0.0010) | ||
| Alternative | 24 | 0.003∗∗ | 0.007∗∗∗ | 0.010∗∗∗ | 0.006∗∗∗ |
| (0.0005) | (0.0007) | (0.0005) | (0.0007) | ||
| ∗p0.05; ∗∗p0.01; ∗∗∗p0.001 | |||||
Appendix D Relationship between awesome and atypical projects.
| Estimated coefficient | |
| Variables of interest | |
| Is awesome | 0.028∗∗∗ |
| Controls | |
| Org_owned | 0.017∗∗∗ |
| (log) | 0.001 |
| (log) | 0.017∗∗∗ |
| (log) | 0.023∗∗∗ |
| Fixed effect | |
| Observations | 37,451 |
| Adjusted R2 | 0.059 |
| ∗p0.05; ∗∗p0.01; ∗∗∗p0.001 |
In Section 3.6, we present that the “awesome” projects on average are more atypical per our measurement than other projects in our sample, and we further validate this result with a regression analysis in Table 9, where we show that this relationship is not driven by confounding variables such as the project size and team size.
Appendix E Real-world examples of projects with varying level of tie diversity
In Section 3.5, we present graphs depicting the networks of projects characterized by varying levels of average diversity and relative weak tie diversity. To concretize those abstract diagrams, we also provide real-world examples in Figure 7.